OpenStudy (coconut):

Find the solution to the equation 64^(x – 3) =4^2x.

OpenStudy (anonymous):

\[64^{(x-3)} = 4^{2x}\] \[4^{3(x-3)} = 4^{2x}\] so 3x-3 = 2x 3x-2x = 3 x = 3

OpenStudy (coconut):

what method di you use

OpenStudy (anonymous):

i just made the base of LHS =RHS then the power of LHS should be equal to RHS

OpenStudy (anonymous):

Um, 3x-9=2x-> Answer is 9